OpenStudy (anonymous):

x^(3/4)=27 How do I solve this equation?

OpenStudy (anonymous):

x=81 Raise each side of the equation to the 4/3 power to eliminate the fractional exponent on the left-hand side. Then solve for X

OpenStudy (jdoe0001):

\(\bf \large x^{\frac{3}{4}}=27 \implies \sqrt[4]{x^3} = 27\\ \textit{raise both sides by 4}\\ \left(\sqrt[4]{x^3}\right)^4 = (27)^4 \implies x^3 = (27)^4\) as suggested by samsterz